Which best describes the given compound propositions [p ∧ (q v r)] <-> [(p ∧ q) v (p ∧ r)]

1
Expert's answer
2021-12-15T16:41:04-0500

Taking into account that "\alpha\leftrightarrow\alpha\equiv T" for any formula "\alpha" and


"[p ∧ (q \lor r)] \leftrightarrow [(p ∧ q) \lor (p ∧ r)]\equiv [p ∧ (q \lor r)] \leftrightarrow [p ∧ (q \lor r)] \equiv T,"


we conclude that this formula is tautology and the logic constant "T" the best describes the given compound proposition.